OpenStudy (anonymous):

Simplify this rational expression:

OpenStudy (anonymous):

\[\frac{ x ^{3}+1 }{ x ^{3}-x ^{2}+x}\times \frac{ 6x }{ -54x-54 }\]

OpenStudy (anonymous):

@Hero

hero (hero):

I usually prefer rational equations, but I guess we can do this one

hero (hero):

Hint: First thing to notice is that in the first fraction, the numerator is a sum of cubes and the denominator, x is common to each term.

hero (hero):

The sum of cubes formula is \((a^3 + b^3) = (a+b)(a^2 - 2ab + b^2)\)

hero (hero):

Also, just for clarity, the numerator of the first fraction can be re-written as \(x^3 + 1^3\) since \(1^3 = 1\)

OpenStudy (anonymous):

I don't understand the whole cubes formula.

hero (hero):

It's just a formula. Keep that in mind. You need to replace the left side of that formula with the right side in your fraction.

OpenStudy (anonymous):

How?

hero (hero):

It's not really that difficult \((a^3 + b^3) = (a+b)(a^2 - 2ab + b^2)\) \((x^3 + 1^3) = (x + 1)(x^2 - 2x + 1)\) Keep in mind that: \(2x \dot\ 1 = 2x\) \(1^2 = 1\)

hero (hero):

It's just an annoying formula. I'm sorry that you do not understand

hero (hero):

But anyway, if you can hold off your confusion for a bit longer, you'll understand that the numerator and denominator would look like this afterwards: \[\large \frac{ (x + 1)(x^2 - 2x + 1)}{ x(x^2 - x + 1)} \times\ \frac{ 6x }{ -54x-54 }\]

hero (hero):

The sum of cubes formula is a factoring method.

hero (hero):

Just a "not so obvious" one. You have to push the "I believe" button in that regard.

hero (hero):

The sum of cubes formula is like the slope-intercept formula. It's not something you necessarily have to understand. It's just something you memorize and apply when you recognize a sum of cubes such as \(a^3 + b^3\)

OpenStudy (anonymous):

Oh my god, guess what I just realized.

OpenStudy (anonymous):

I am so sorry, and I would hate me if I were you. But, the instructions were actually Multiply as indicated...

hero (hero):

I already know that. It is obvious that you need to multiply the fractions since there is a multiplication symbol between the fractions, but you assume that you can just multiply as simple as that. In this case, we need to simplify the left fraction BEFORE multiplying it with the right fraction.

OpenStudy (anonymous):

Ohh, okay. Well, I know how to factor the right side, and whatnot, I wasn't sure how to factor the left

hero (hero):

The left fraction still needs to be simplified BEFORE you multiply them together.

hero (hero):

The left fraction is not fully simplified yet

OpenStudy (anonymous):

Ohh, okay.

hero (hero):

You should be able to recognize a quadratic expression when you see one and then figure out if it can be factored. You do see quadratic expressions in the numerator and denominator, right?

OpenStudy (anonymous):

Yeah

hero (hero):

Okay, well, let me know the factored forms of each one.

OpenStudy (anonymous):

maybe..\[x ^{2}(x+1)\] for the numerator?

hero (hero):

You're losing me here. You need to show all of your work.

OpenStudy (anonymous):

I guess I'm confused as to how I can factor the left

hero (hero):

I guess I assumed you already knew how to factor

OpenStudy (anonymous):

Well, I do...but not necessarily for the left side..

hero (hero):

@GenGen96

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'm sorry, I honestly don't mean to rush, but there is only this question nd one more and then I have to head off to bed, it's 11:25PM here

hero (hero):

Okay, you said you factored it already so what do you have so far?

OpenStudy (anonymous):

I have: \[\frac{ (x+1)(x ^{2}-+1) }{ x(x ^{2}-x+1) }\times \frac{ 6x }{ -54(x+1) }\]

hero (hero):

From my end, you appear to have forgotten the 2x in the numerator of the first fraction. Is this what you have: \[\large \frac{ (x + 1)(x^2 - 2x + 1)}{ x(x^2 - x + 1)} \times\ \frac{ 6x }{ -54x-54 }\]

OpenStudy (anonymous):

i forgot the x, but why is there a 2x?

hero (hero):

Oh, I see. I wrote down the formula wrong.

OpenStudy (anonymous):

I did forget the x, though, so yes.

hero (hero):

Okay, well after making corrections, this is what you should have: \[\large \frac{ (x + 1)(x^2 - x + 1)}{ x(x^2 - x + 1)} \times\ \frac{ 6x }{ -54(x+1) }\]

hero (hero):

And actually, we can immediately cancel something, can't we?

hero (hero):

Do you see a factor of 1 that we can cancel?

OpenStudy (anonymous):

Ohh, okay.

hero (hero):

Since you think you have it figured out, let me know what the final result is.

OpenStudy (anonymous):

do I cancel the x^2-x+1 too?

hero (hero):

Yup. There are plenty of things that cancel

hero (hero):

In a situation like this, what you want to do is organize everything so that you can keep track of what you need to cancel.

OpenStudy (anonymous):

\[\frac{ 1 }{ x }\times \frac{ 6x }{ -54 }\]?

hero (hero):

\[\large \frac{ 6x(x + 1)(x^2 - x + 1)}{ -54x(x+1)(x^2 - x + 1)} \]

hero (hero):

Now it is more obvious what needs to be canceled

OpenStudy (anonymous):

WHat about that x that was on the outside when everything was factored?

hero (hero):

I included that as well, trust me

hero (hero):

It's just hard for you to see it maybe

OpenStudy (anonymous):

ohhh i see it

OpenStudy (anonymous):

Okay, i got the answer. Thank you! Okay, last question..

hero (hero):

You don't have the answer until you let me know what it is. It shouldn't be very difficult to type at all.

OpenStudy (anonymous):

-1x/9x

hero (hero):

The x's don't cancel?

OpenStudy (anonymous):

Ohhhhh! Okay

OpenStudy (anonymous):

so -1/9!

hero (hero):

lol

hero (hero):

Great job :D

OpenStudy (anonymous):

:D Thank you! I'll start the question in a new thing and then close it! One second!

OpenStudy (anonymous):

Okay, I mentioned you